# Engineering Scientific Assistants using Interactive Structured Induction of Programs

### v1

@@ Abstract
In a lecture in 1965 on “Programming Considered as A Human Activity” [1] Edsger Djikstra initiated what would later be called structured programming. mainly to make programs both human-understandable, and amenable to proofs of correctness. He suggested complex programs should be constructed like mathematicians construct complex proofs: start with some known primitives and libraries (axioms), construct intermediate sub-programs that are verified independently for correctness (lemmas), and then arrive at the program by reuse of the sub-programs (the “proof”).
As is evident from the title of his lecture, Djikstra envisages programs will be written and understood by human programmers. While the possibility of machine-authored programs was envisaged as early as Turing’s seminal 1950 paper [2], it is probably the mid-1980s when machine learning (ML) techniques were producing sufficiently complex programs to raise concerns on both their (human-) understandability and correctness (this is even true of ML techniques that constructed output in symbolic representations: see [3]).

### v2

@@ Abstract
In The Mythical Man-Month, Fred Brooks famously identified software as perhaps the most complex of human creations [1], although the idea of programs as exceedingly complex artifacts had already been raised by Herb Simon. But well before both Brooks and Simon—focussing on the artisan rather than the artifact—Dijkstra stated: “As the power of available computing increases, the problems tackled by programmers have grown ever more complex, and programming itself has become a most challenging intellectual endeavour.” [2]. Today, practising software engineers may be even more sympathetic to Dijkstra’s dilemma than their counterparts of five decades ago. To date, algorithm design and program construction have remained largely resistant to significant automation. However, the advent of foundation models trained on billions of lines of code may be poised to change this situation. Large language models (LLMs) are now being used to generate code for a wide range of tasks, from routine programming (for example, form-filling and basic web design) to more open-ended activities, such as automating aspects of scientific research [3, 4, 5, 6]. Unsurprisingly, as LLMs become increasingly capable at code generation, Dijkstra resurfaces: software engineers—now armed with LLMs—are expected to construct software for problems where human prior knowledge and intuition may be limited or even absent. We are at an early stage of understanding how best to use tools like LLMs for program construction.